# Coupled Microelectromechanical Drum Resonators for Reservoir Computing via Sideband Pumped Phonon-Cavity Dynamics

### v1

@@ 2.3 Reservoir computing implementation in the phonon-cavity scheme
The values of the constants $\alpha_{1}$, $\alpha_{2}$ and offsetoffset are fixed by programmed MATLAB codes to avoid the saturation of the FPGA card, $x_{0}$ denotes the preset offset in the measurement. These values define the modulation range of the pump amplitude that can generate the nonlinearity required for implementing reservoir computing.
@@ 2.4 The Parity Benchmark
$P_{n}(t)=\prod_{i=0}^{n-1}u[t-(i+\delta)\tau].$ (4)
In the experiment, we set $\delta=0$. The objective is to train the output weights of the system to provide an output which can be as close as possible to the target output $P_{n}$. For $n>1$, the the parity function is nonlinear separable and the task requires the reservoir to store information about a nonlinear transform of previous inputs [Sylvestre].
@@
Figure 3: Reservoir Computing based on the sideband pumping mechanism while probing the SiN drum with $\Omega_{d}/(2\pi)=12.26$ MHz, $V_{d}=6$ mV, $\Omega_{p}/(2\pi)=18.38$ MHz and $V_{dc}=4$ V. (a) The detected mechanical displacement of the probed SiN membrane as a function of the amplitude of the blue pump signal. The curve is divided into 4 modulation windows $M1$, $M2$, $M3$, and $M4$ showcasing different profile variations depending on the interval of the blue pump voltage. (b) Comparison between the predicted output by our reservoir computing scheme (black lines) and the target ideal output of the parity benchmark (red lines) for the first 7 orders of the parity benchmark in the case of the modulation window $M4$. (c) Success rates obtained for the first 4 orders of the Parity Benchmark corresponding to each of the 4 modulation windows.
@@
Fig.3 (b) shows one example of the comparison between the target values $P_{n}(t)$ and the predicted values by the neural network, which are obtained in the modulation window $M4$, with $\theta=100$ $\mu s$ and a period of $\tau=N_{mask}\theta=40$ ms. The values of the constants $\alpha_{1}$, $\alpha_{2}$ and offsetoffset that modulate the electrostatic pump force were set respectively to 0.55, 0.34 and 0.69. The success rates show near-perfect results for the first-order $P_{1}$ to the third-order $P_{3}$ of the parity benchmark, then the curve becomes more noisy for the $P_{7}$. This decrease of the success rate is linked to the requirement of a higher memory capacity from our device.
@@
Figure 3 (c) presents the reservoir computing results for the different modulation windows described above. The best results are obtained for $M4$ where the characterization curve is strongly nonlinear and where the influence of the interferences are the most dominant. Consequently, the $M3$ modulation window yields the second-best performance due to its nonlinear response profile. However, as for the $M1$ window, very poor results are achieved reaching approximately 50 % starting from $P_{2}$. This is due to the absence of nonlinearity in this modulation window and low signal to noise ratio. Lastly, the window $M2$ in which the force is modulated between 0.5 mV and 175 mV shows also poor results reaching around 60 % at the second-order. In this modulation range, the linear plateau spans a much wider interval than the nonlinear region, reducing the likelihood that the modulated pump amplitude enters the $V_{p}$ range that generates nonlinear behavior. Consequently, the success rates decrease rapidly, indicating poor reservoir computing performance for this modulation window.
@@
Figure 4: reservoir computing based on the sideband pumping mechanism while probing the Al drum with $\Omega_{d}/(2\pi)=6.12$ MHz, $V_{d}=80$ mV, $\Omega_{p}/(2\pi)=18.38$ MHz and $V_{dc}=4$ V. (a) The detected mechanical displacements of the probed Al membrane as a function of the variation of the amplitude of the blue pump signal. The curve is divided into three modulation windows of the pump amplitudes the $M1$, $M2$ and $M3$ where the detected probe tone exhibits different variations. (b) Success rates for the first four orders of the parity benchmark (from $P_{1}$ to $P_{4}$) for the three different modulation windows. (c) The values of the obtained NMSE in the case of NARMA2 and NARMA10 corresponding to each modulation window
@@
Probing the Al drum resonator We use the same probing method described above for the phonon cavity to obtain the nonlinear amplification curve of the probe tone around the $\Omega_{Al}$, as shown in Fig. 4 (a). This curve was subdivided into 3 different modulation windows: $M1$ marked by a linear plateau fluctuating around a minimum value of 0.2 mV, $M2$ where the curve exhibits a nonlinear amplification up to a maximum of 0.24 mV accompanied by a fast drop and $M3$ that combines the $M2$ window with the sudden jump to a displacement value of 0.34 mV. The reservoir computing studies with these 3 modulation windows were conducted with values of the force modulation parameters $\alpha_{1}$, $\alpha_{2}$ and offset respectively 0.56, 0.33 and 0.71. Promising success rates were obtained in the cases of the windows $M2$ and $M3$ due to the nonlinear variation of the characterization curve in each window. Near perfect results were obtained for both $P_{1}$ and $P_{2}$ in both cases and then it decreases slowly to reach around 70% for the $P_{4}$. As for $M1$, the success rates deteriorated quickly reaching an unsatisfactory 50% starting from the second-order $P_{2}$ of the parity benchmark due to the absence of nonlinear variations within this modulation range of the pump force.
@@ 2.5 The NARMA Benchmark
The Normalized Auto-Regressive Moving Average (NARMA) benchmark is widely used to evaluate nonlinear processing and fading memory in reservoir computing systems. The key feature of the NARMA benchmark is that the current output is generated through a nonlinear combination of many past inputs and outputs. As a result, it requires the reservoir that combines nonlinear processing with long fading memory [barazani2020microfabricated]. A generalized version of its input-output relationship for this benchmark is given by:
$\hat{y}_{n}(t+1)=0.3\hat{y}_{n}(t)+0.05\hat{y}_{n}(t)\sum_{i=0}^{n-1}\hat{y}_{n}(t-i)+1.5u(t)u(t-n+1)+0.1$ (5)
where $n$ is a time-lag parameter. The input $u(t)$ of the system consists of scalar random numbers, drawn from a uniform distribution over the interval $[0,0.5]$. The $\hat{y}_{n}$ is the output target. In order to evaluate the performance of the system based on the NARMA benchmark, we compute the Normalised Mean Squared Error (NMSE). Assuming that the target output vector is $\hat{y}$ and the predicted output vector by our reservoir computing is $y$, we adopt the following definition of the NMSE [NARMA10, guo2024mems]:
[…]
As demonstrated in previous studies on the parity benchmark, the initial step in designing a reservoir computing framework is to determine the optimal pump force modulation window. Here, we evaluated the system’s NARMA performance across various modulation windows by probing the Al drum resonator via a sideband pumping scheme, as presented in Fig. 4(a). The resulting measurements for each modulation window are shown in Fig. 4(c). The optimal NMSE values were achieved at modulation window $M3$, coinciding with the nonlinear jump where phonon interference effects become dominant. This observation aligns with our previous results from the parity benchmark tests.
@@
By selecting an optimal modulation window that maximizes the system’s nonlinear response, we evaluated various modulation schemes using the NARMA benchmark. The first approach leverages the Duffing nonlinearity, the predominant nonlinear characteristic of microelectromechanical systems, which has been successfully employed for reservoir computing in several prior studies [Sylvestre, zheng2021parameters]. In our double-drum architecture, we exploit the Duffing properties of a SiN circular resonator. Due to its specific mechanical geometry, this design typically yields a higher nonlinearity coefficient than conventional doubly-clamped beams [cattiaux, venkatachalam2023effects]. The second and third ones are based on the sideband pumping creating nonlinear amplification mechanism while probing respectively the SiN phonon-cavity and Al resonators. The experimental findings are presented in Fig. 5 and details regarding the training and testing parameters can be found in table 1.
@@
In the case of Duffing nonlinearity (Fig. 5(a)), the NMSE increases gradually from 0.2 and converges to approximately 0.6 at the tenth-order NARMA benchmark. The same trend is observed in the sideband pumping experiment while probing SiN (Fig. 5(b)) and Al drums (Fig. 5(c)). It is noticeable that the NMSE increases slightly while probing the SiN drum as it converges to approximately 0.7 and it reaches 0.8 for the NARMA10 in the case of probing the Al drum.
@@ 3 Discussion
We compare the performance of this MHz-range double-drum resonator with previously reported reservoir computing implementations using single MEMS resonators operating at kHz frequencies. For the parity benchmark, the success rates achieved in this double-drum resonator with the sideband-pumping approach are slightly lower than those obtained using the Duffing nonlinearity alone. These results are comparable to those reported for MEMS accelerometer based reservoirs [barazani2020microfabricated], but remain below the performance achieved with doubly clamped beam MEMS implementations [Sylvestre]. Regarding the NARMA benchmark, our NMSE results do not yet match the good performance of the kHz-range MEMS, particularly those utilizing stiffness modulation techniques [guo2024mems, Sun2021, zheng2021parameters]. We attribute this behavior to additional noise associated with the fast-decaying Al mode, which influences the overall reservoir computing performance of the double-drum system. This work represents the first demonstration of reservoir computing using coupled MEMS resonators operating in the MHz frequency regime. However, the high-frequency operation imposes stricter constraints on the fading memory compared to the kHz systems. Further optimizations of both training algorithm and device design are required to improve performance. One promising direction is to enhance the reservoir computing algorithm at the software level, for example by optimizing post-processing strategies to extend the effective long term memory capacity [guo2024mems]. In parallel, the mechanical properties of the Al drum could be improved through optimization of the nanofabrication process, such as reducing clamping losses or increasing intrinsic stress to enhance the quality factor and thereby suppress noise and further reducing the required driving forces.
This work presents a novel reservoir computing scheme based on a double-drum resonator system that leverages concepts from phonon-cavity electromechanics. In the two-tone scheme, nonlinear interference effects are induced by applying a pump tone at the blue sideband of the phonon cavity while probing one of the coupled resonators. By exploiting this effect and modulating the pump amplitude, we experimentally demonstrate time delay loop based reservoir computing using two capacitively coupled drum resonators. The pump tone acts as a coherent data bus, generating phonons in the unprobed mode, which are subsequently projected back to the probed mode forming an interference effect. This sideband pumping method allows one of the coupled mode to function as a dedicated sensor, while the detected signals are transduced into mechanical properties, for instance the variations of phonon numbers or resonance frequency, and coherently transferred to the second drum for reservoir computing. This separation of sensing and processing within a single electromechanical platform advances the development of compact MEMS systems that integrate sensing and computing functionalities. In addition, we would like to emphasize that the proposed reservoir computing scheme based on the sideband pumping technique is not limited to our double-drum resonator architecture and does not require the resonance frequencies of the two coupled resonators to be closely matched. It can be straightforwardly extended to other multimode coupling platforms, including mechanically coupled resonator arrays and optomechanical systems.
@@ 4 Materials and methods
Implementation of measuring equipment: The microwave source used for microwave interferometry measurements is an Rohde &\& Schwarz SMB100B signal generator, which provides ultra-low phase noise. The microwave power shining to the double-drum through transmission line is around -5 dBm. A Yokogawa GS200 source is used to supply the $DC$ bias voltage $V_{dc}$. The $AC$ drive signals are generated using a Keysight Technologies 33622A arbitrary waveform generator for the pump tone and a Zurich Instruments UHFLI 600 MHz lock-in amplifier for the probe tone. After frequency down-conversion of the microwave signal to the MHz range, the mechanical displacement excited by the probe tone is detected by the lock-in amplifier as a voltage signal. This signal is then fed back to the input of a AMD Spartan-7 FPGA (XC7S100) to create a time delay $\tau$. It is then added with the the masked data (mask $m(t)$ onto the input data $u(t)$) for the output of the FPGA. An AD835 analog multiplier is used to multiply the incident pump signal with the FPGA output, in order to modulate the pump force applied to the double-drum resonator.
Training: The displacement vector, created from the force amplitude modulation, will be linearly combined and give the final output $y(t)$ based on the formula
@@
The Z is the data displacement matrix consisting of $x(t)$ and $\lambda$ is the regularization parameter. When the training of the weight vector $W$ ends, we will test the performance of our system for different inputs [Brunner_ridge_regression, Sylvestre]. All computing processes are based on 2000 input bits. The first 100 inputs are discarded to eliminate offset effects from previous measurements. The subsequent 1500 inputs are used for training, and the remaining 400 inputs are used for prediction.
@@
X.Z. conceived the experiments and supervised the project. L.F. and X.Z. fabricated the samples in the clean room. P.L. and X.Z. developed the experimental setup. T.F., L.F., and X.Z. conducted the experiments and analyzed the results. All authors reviewed the manuscript and contributed to the discussions.
Additional information
5 Supplementary Information
5.1 Reservoir Computing based on the amplitude modulation of the force in the Duffing regime
Micro/nano-mechanical resonators have intrinsic nonlinear properties, coming from material properties and geometries. One well-known phenomenon is the Duffing behavior, in which the frequency response of MEMS under strong driving no longer follows the simple Lorentzian form. The motion equation of the resonator becomes :
$\ddot{x}+\gamma_{m}\dot{x}+\Omega_{m}^{2}x+\beta x^{3}=\frac{F_{d}}{m_{eff}}\cos({\omega_{d}t})$ (9)
where $\gamma_{m}$ is the damping coefficient, $\Omega_{m}$ is the resonance frequency of the resonator, $\beta$ is the Duffing nonlinearity coefficient normalized by the effective mass $m_{eff}$ and $F_{d}$ is the amplitude of the electrostatic force. For a circular membrane drum mechanical resonator (e.g. SiN drum or Al drum), the Duffing nonlinearity is primarily geometric in origin [cattiaux]. In this work, we excite only the fundamental mode of the SiN drum in the Duffing nonlinear regime. No Duffing nonlinearity was observed in the Al drum resonator.
Figure 6: Reservoir Computing based on Duffing nonlinearity. (a) The mechanical response of the SiN resonator as a function of the frequency for $V_{d}=15$ mV, attenuated by 20 dB. The red curve corresponds to the reverse frequency sweep and the black curve corresponds to the forward sweep. (b) The variation of the mechanical response of the resonator as a function of the drive voltage at the frequency $\Omega_{d}/(2\pi)=12.257$ MHz. The curve is divided into 3 modulation windows of the electrostatic force applied: $M1$, $M2$ and $M3$. (c) The results of reservoir computing based on the modulation of the electrostatic drive force in the Duffing nonlinear regime for the 3 different modulations windows. (d) Comparison between the target output in red and the predicted output after training the output weights in black, which are obtained in the modulation window $M3$.
Figure 6 (a) shows the characteristic Duffing nonlinear frequency response of the SiN drum resonator when driven by a relatively large electrostatic force. For $\beta>0$, strong driving causes the resonance frequency to shift toward higher values, and the frequency response exhibits hysteretic behavior between forward (black curve) and backward (red curve) frequency sweeps. The hysteresis window is bounded by two saddle-node bifurcation points, within which stochastic switching between bistable states can arrive [venkatachalam2023effects]. For reservoir computing experiments, the operating frequency is chosen outside the bistable region at $\Omega_{d}/(2\pi)=12.257$ MHz, as indicated by the dashed line in Fig. 6 (a). By fixing the drive frequency at this value, we measure the mechanical response as a function of the driving force amplitude $V_{d}$, as shown in Fig. 6 (b). Based on this response, three different driving-force modulation windows are selected for evaluating the reservoir computing performance. The $M3$ represents the whole window with variations of the drive amplitude between 0 and 16 mV where the curve shows a clear nonlinear profile. Similarly, the $M1$ showcases nonlinear variations, it is proper to drive voltage amplitudes less than 9 mV thus limiting the electrostatic drive force to smaller values. Lastly, in the $M2$ the variation of the mechanical response of the SiN membrane as a function of the drive amplitude is almost linear.
When the $M3$ modulation window is used, near-perfect success rates are achieved from the first-order ($P_{1}$) to the fourth-order ($P_{4}$) parity benchmarks, owing to the strong nonlinear variations present in this regime. As for the window $M1$, the results surpass a value of 90 % for $P_{1}$ and $P_{2}$ indicating successful predictions by our system. However, they drop to approximately 58 % starting from $P_{3}$. Moreover, for $M2$ where there is an absence of nonlinearity, the results are less important reaching 70 % for $P_{2}$ and then dropping quickly to approximately 59 % starting from $P3$.
The same conclusion can be taken, the optimal results are obtained in the modulation window that exhibits the most amount of nonlinearity in the amplitude-amplitude curve.
5.2 Two-tone driving scheme
The coupled electromechanical resonators act as capacitor $C_{g}(X_{1},X_{2})$ consisting of 2 moving membranes. The mechanical displacement amplitudes are expressed as $X_{1}(t)$ and $X_{2}(t)$ respectively for each of the SiN and Al drum resonators resonating at frequencies $\Omega_{1}$ and $\Omega_{2}$. The motion equation of this double-drum resonator is described by
$\ddot{X_{2}}+\gamma_{2}\dot{X_{2}}+\Omega_{2}^{2}X_{2}$ $=\frac{V_{ac}V_{dc}}{m_{2}d}C_{go}\left[-1+2\frac{({X_{2}}-{X_{1}})}{d}\right].$
In the blue sideband pumping scheme, $\Omega_{p}$ = $\Omega_{1}+\Omega_{2}+\Delta$, we probe the Al drum at the frequency $\Omega_{d}=\Omega_{2}+\delta$. The mechanical displacement $x_{1}$ and $x_{2}$ corresponding respectively to the unprobed and probed resonators are established as follows:
$x_{1}$ $=-\frac{f_{p}}{2m_{1}\Omega_{1}}\frac{x_{2}^{*}}{d}\chi_{1},$ (11)
$x_{2}$ $=-\frac{f_{d}}{2m_{2}\Omega_{2}}\frac{1}{\frac{1}{\chi_{2}}-\frac{|f_{p}|^{2}\chi_{1}^{*}}{4m_{1}m_{2}d^{2}\Omega_{1}\Omega_{2}}}.$
The $f_{p}=\frac{C_{g0}V_{dc}\mu_{p}}{d}$ and $f_{d}=\frac{C_{g0}V_{dc}\mu_{d}}{d}$ are the complex amplitudes of the pumping and the driving forces. The definition of the mechanical susceptibilities become:
$\chi_{1}$ $=\frac{1}{\delta-\Delta-i\frac{\gamma_{1}}{2}},$ (12)
$\chi_{2}$ $=\frac{1}{-\delta-i\frac{\gamma_{2}}{2}}.$
All derivation details have been reported in our previous work [alok].
Figure 7: (a) Frequency responses of Al drum measured $V_{d}=80$ mV, when the phonon-cavity is pumped at its blue sideband at frequency $\Omega_{p}=\Omega_{1}+\Omega_{2}=2\pi\times 18.3$ MHz and a pump amplitude of $V_{p}=0.11$ V. (b) The 2D plots of the frequency responses of the SiN drum resonator corresponding to different blue sideband pump amplitudes.
In the two-tone scheme, the interference window is primarily determined by the bandwidths of the two coupled modes, as reflected in the denominators of Eqs. 11 and 12. For the case of probing Al drum, the bandwidth of SiN $\gamma_{SiN}$ is much smaller than that of Al drum, $\gamma_{SiN}\ll\gamma_{Al}$. As a result, the interference effect induced by blue-sideband pumping appears as a narrow peak within the broader bandwidth of the Al drum [alok]. The pump frequency is set to $\Omega_{SiN}$ + $\Omega_{Al}$ + $\delta$. When the swept probe frequency around the Al drum resonance reaches $\Omega_{Al}$ + $\delta$, this interference effect becomes clearly observable, as shown in Fig.7(a).
In the sideband pumping scheme, the resonance frequency of the coupled drum resonator will shift because the pump amplitude modifies the susceptibility of the electromechanical resonator [alok]. It is well-known in the optomechanical system, so called the optical spring effects. In the blue sideband pump scheme, we sweep frequency responses of the SiN drum corresponding to different pump amplitudes. As present in the Fig 7(b), the maximal resonance frequency $\Omega_{SiN}$ decreases from $2\pi\times 12.272$ MHz to $2\pi\times 12.265$ MHz when the pump amplitude increases from 0 V to $125$ mV. This $2\pi\times 7$ kHz shift is still within the bandwidth of the Al membrane ($\gamma_{2}\approx 2\pi\times 25.9$ kHz), however, it is much larger than that of the SiN membrane ($\gamma_{1}\approx 2\pi\times 944$ Hz). In the two-tone scheme, the modulations of the pump amplitudes therefore can bring the probe tone into or out of the interference window, thereby forming a controllable nonlinear response.

### v2

@@ 2.3 Reservoir computing implementation in the phonon-cavity scheme
The values of the constants $\alpha_{1}$, $\alpha_{2}$ and offsetoffset are fixed by programmed MATLAB codes to avoid the saturation of the FPGA card, $x_{0}$ denotes the preset offset in the measurement. These values define the modulation range of the pump amplitude that can generate the nonlinearity required for implementing reservoir computing. In the phonon-cavity two-tone scheme, the masked input data $u(t)m(t,\theta)$ are applied to the pump force $f_{p}$, which transfers the information to the mechanical displacements of the two coupled drum resonators. Figure 2 (e) present the schematic of this concept. In principle, based on Eq. 1, the displacement of either drum resonator could be used for the training process. However, in our experiment we choose to read out the probed resonator in order to have the better signal to noise ratio.
@@ 2.4 The Parity Benchmark
$P_{n}(t)=\prod_{i=0}^{n-1}u[t-i\cdot\tau].$ (4)
The objective is to train the output weights of the system to provide an output which can be as close as possible to the target output $P_{n}$. For $n>1$, the parity function is nonlinear separable and the task requires the reservoir to store information about a nonlinear transform of previous inputs [11].
@@
Figure 3: Reservoir computing based on the sideband pumping mechanism while probing the SiN drum with $\Omega_{d}/(2\pi)=12.26$ MHz, $V_{d}=6$ mV attenuated by 20 dB, $\Omega_{p}/(2\pi)=18.38$ MHz and $V_{dc}=4$ V. (a) The detected mechanical displacement of the probed SiN membrane as a function of the amplitude of the blue pump signal. The curve is divided into 4 modulation windows $M1$, $M2$, $M3$, and $M4$ showcasing different profile variations depending on the interval of the blue pump voltage. (b) Comparison between the predicted output by our reservoir computing scheme (black lines) and the target ideal output of the parity benchmark (red lines) for the first 7 orders of the parity benchmark in the case of the modulation window $M4$. (c) Success rates obtained for the first 4 orders of the parity benchmark corresponding to each of the 4 modulation windows.
@@
Fig.3 (b) shows one example of the comparison between the target values $P_{n}(t)$ and the predicted values by the neural network, which are obtained in the modulation window $M4$, with $\theta=100$ $\mu s$ and a period of $\tau=N_{mask}\theta=40$ ms. We set the constants $\alpha_{1}$, $\alpha_{2}$ and offsetoffset to 0.55, 0.34 and 0.69, respectively, which are obtained from an optimization process designed to achieve the best success rates while avoiding saturation of the FPGA card. The success rates show near-perfect results for the first-order $P_{1}$ to the third-order $P_{3}$ of the parity benchmark, then the curve becomes more noisy for the $P_{7}$. This decrease in the success rate can be attributed to the lack of higher memory capacity in our device.
@@
Figure 3 (c) presents the reservoir computing results for the different modulation windows described above. The best results are obtained for $M4$ where the characterization curve is strongly nonlinear and where the influence of the interferences are the most dominant. Consequently, the $M3$ modulation window yields the second-best performance due to its nonlinear response profile. However, as for the $M1$ window, very poor results are achieved reaching approximately 50 % starting from $P_{2}$. This is due to the absence of nonlinearity in this modulation window and low signal to noise ratio. Lastly, the window $M2$ in which the force is modulated between 0.5 mV and 175 mV shows also poor results reaching around 60 % at the second-order. In this modulation range, the linear plateau spans a much wider interval than the nonlinear region, reducing the likelihood that the modulated pump amplitude enters the $V_{p}$ range that generates nonlinear behavior. Consequently, the success rates decrease rapidly, indicating poor reservoir computing performance for this pump tone modulation window.
@@
Figure 4: Reservoir computing based on the sideband pumping mechanism while probing the Al drum with $\Omega_{d}/(2\pi)=6.12$ MHz, $V_{d}=80$ mV (before a 20 dB attenuation), $\Omega_{p}/(2\pi)=18.38$ MHz and $V_{dc}=4$ V. (a) The detected mechanical displacements of the probed Al membrane as a function of the variation of the amplitude of the blue pump signal. The curve is divided into three modulation windows of the pump amplitudes the $M1$, $M2$ and $M3$ where the detected probe tone exhibits different variations. (b) Success rates for the first four orders of the parity benchmark (from $P_{1}$ to $P_{4}$) for the three different modulation windows. (c) The values of the obtained NMSE in the case of testing NARMA2 and NARMA10 tasks, corresponding to each pump force modulation window shown in the (a).
@@
Probing the Al drum resonator We use the same probing method described above for the phonon cavity to obtain the nonlinear amplification curve of the probe tone around the $\Omega_{Al}$, as shown in Fig. 4 (a). This curve was subdivided into 3 different modulation windows: $M1$ marked by a linear plateau fluctuating around a minimum value of 0.2 mV, $M2$ where the curve exhibits a nonlinear amplification up to a maximum of 0.24 mV accompanied by a fast drop and $M3$ that combines the $M2$ window with the sudden jump to a displacement value of 0.34 mV. The reservoir computing studies with these 3 modulation windows were conducted with values of the force modulation parameters $\alpha_{1}$, $\alpha_{2}$ and offset respectively 0.56, 0.33 and 0.71, in order to have the best success rate. Promising success rates were obtained in the cases of the windows $M2$ and $M3$ due to the nonlinear variation of the characterization curve in each window. Near perfect results were obtained for both $P_{1}$ and $P_{2}$ in both cases and then it decreases slowly to reach around 70% for the $P_{4}$. As for $M1$, the success rates deteriorated quickly reaching an unsatisfactory 50% starting from the second-order $P_{2}$ of the parity benchmark due to the absence of nonlinear variations within this modulation range of the pump force.
@@ 2.5 The NARMA Benchmark
The NARMA benchmark is widely used to evaluate nonlinear processing and fading memory in reservoir computing systems. The key feature of the NARMA benchmark is that the current output is generated through a nonlinear combination of many past inputs and outputs. As a result, it requires the reservoir that combines nonlinear processing with long fading memory [5]. A generalized version of its input-output relationship for this benchmark is given by:
$\hat{y}_{n}(k+1)=0.3\hat{y}_{n}(k)+0.05\hat{y}_{n}(k)\sum_{i=0}^{n-1}\hat{y}_{n}(k-i)+1.5u(k)u(k-n+1)+0.1$ (5)
where $n$ is a time-lag parameter, $k=\frac{t}{\tau}$ is the timestep. The dimension of the input and output vectors corresponds to a total of $N$ data divided between $N_{tr}$ for the training phase and $N_{tst}$ that are utilized to test and evaluate the performance of the reservoir computing experiment. The input $u(k)$ of the system consists of scalar random numbers, drawn from a uniform distribution over the interval $[0,0.5]$. The $\hat{y}_{n}$ is the output target. In order to evaluate the performance of the system based on the NARMA benchmark, we compute the Normalised Mean Squared Error (NMSE). Assuming that the target output vector is $\hat{y}$ and the predicted output vector by our reservoir computing is $y$, we adopt the following definition of the NMSE [38, 17]:
$\begin{split}MSE(\hat{y}_{n},y_{n})=\frac{1}{N_{tst}}\sum_{i=1}^{N_{tst}}(\hat{y}_{n}(i)-y_{n}(i))^{2}\\ NMSE(\hat{y}_{n},y_{n})=\frac{MSE(\hat{y}_{n},y_{n})}{MSE(\hat{y}_{n},mean(\hat{y}_{n}))}.\end{split}$ (6)
As demonstrated in previous studies on the parity benchmark, the initial step in designing a reservoir computing framework is to find the optimal pump force modulation window. Here, we evaluated the system’s NARMA performance across various modulation windows by probing the Al drum resonator, via a sideband pumping scheme, as presented in Fig. 4(a). The resulting measurements for each modulation window are shown in Fig. 4(c). The optimal NMSE values were achieved at the modulation window $M3$, coinciding with the nonlinear jump where phonon interference effects become dominant. This observation is consistent with our previous parity benchmark results, indicating that more complex nonlinear behavior benefits for mapping the input into higher-dimensional nonlinear spaces, thereby leading to the improved computing performance.
@@
By selecting an optimal modulation window that maximizes the system’s nonlinear response, we evaluated various modulation schemes using the NARMA benchmark. The first approach leverages the Duffing nonlinearity, the predominant nonlinear characteristic of microelectromechanical systems, which has been successfully employed for reservoir computing in several prior studies [11, 42]. In our double-drum architecture, we exploit the Duffing properties of the SiN drum resonator. More details can be found in the SI part. Due to its circular geometry, the drum design yields a higher nonlinearity coefficient than conventional doubly-clamped beams [7, 35]. The second and third methods are based on sideband pumping, which induces nonlinear phonon-transfer dynamics between two coupled vibrational modes, while probing the SiN phonon cavity and the Al resonator, respectively. The experimental findings are presented in Fig. 5 and details regarding the training and testing parameters can be found in table 1. In the case of Duffing nonlinearity (Fig. 5(a)), the NMSE increases gradually from 0.2 and converges to approximately 0.6 at the tenth-order NARMA benchmark. The same trend is observed in the sideband pumping experiment while probing SiN (Fig. 5(b)) and Al drums (Fig. 5(c)). It is noticeable that the NMSE increases slightly while probing the SiN drum as it converges to approximately 0.7 and it reaches 0.8 for the NARMA10 in the case of probing the Al drum.
@@
We compare the performance of this MHz range double-drum resonator with previously reported reservoir computing implementations using single MEMS resonators operating at kHz frequencies. For the parity benchmark, the success rates achieved in this double-drum resonator with the sideband pumping approach are slightly lower than those obtained using the Duffing nonlinearity alone. These results are comparable to those reported for MEMS accelerometer based reservoirs [5], but remain below the performance achieved with doubly clamped beam MEMS implementations [11]. Regarding the NARMA benchmark, our NMSE results do not yet match the good performance of the kHz range MEMS, particularly those utilizing stiffness modulation techniques [17, 32, 42]. It should be noted that operation in the MHz frequency regime inherently imposes stricter constraints on fading memory compared to kHz MEMS reservoirs, which naturally affects performance on long-memory benchmarks such as NARMA. We attribute the remaining performance gap primarily to additional noise associated with the fast decaying Al mode, which influences the overall reservoir computing performance of the double-drum system. Further optimizations of both training algorithm and device design are required to improve performance. One promising direction is to enhance the reservoir computing algorithm at the software level, for example by optimizing post-processing strategies to extend the effective long term memory capacity [17]. In parallel, the mechanical properties of the Al drum could be improved through optimization of the nanofabrication process, such as reducing clamping losses or increasing intrinsic stress to enhance the quality factor and thereby suppress noise and further reducing the required driving forces.
@@ 3 Discussion
In summary, this work presents a novel reservoir computing scheme based on a double-drum resonator system that leverages concepts from phonon-cavity electromechanics. In the two-tone scheme, controllable nonlinear phonon transfer dynamics between two coupled resonators are induced by applying a pump tone at the blue sideband of the phonon-cavity while probing one of the coupled resonators. The pump tone acts as a coherent data bus, enabling the masked input data to be mapped onto both coupled drums for reservoir computing. Although further efforts are still required to improve reservoir computing performance, this sideband pumping approach provides a way for extending a single MEMS based reservoir computing to multimode systems. In this double-drum design, one of the coupled modes to function as a dedicated sensor, while the detected signals are transduced into mechanical properties, for instance the variations of phonon numbers or resonance frequency, and are coherently transferred to the second resonator for reservoir computing. This separation of sensing and processing within a single electromechanical platform advances the development of compact MEMS systems that integrate sensing and computing functionalities. Finally, we would like to emphasize that the proposed reservoir computing scheme based on the sideband pumping technique is not limited to our double-drum resonator architecture and does not require the resonance frequencies of the two coupled resonators to be closely matched. It can be straightforwardly extended to other multimode coupling platforms, including mechanically coupled resonator arrays and optomechanical systems.
@@ 4 Materials and methods
Implementation of measuring equipment: The microwave source used for microwave interferometry measurements is an Rohde &\& Schwarz SMB100B signal generator, which provides ultra-low phase noise. The microwave power shining to the double-drum through transmission line is around -5 dBm. A Yokogawa GS200 source is used to supply the $DC$ bias voltage $V_{dc}$. The $AC$ drive signals are generated using a Keysight Technologies 33600A arbitrary waveform generator for the pump tone and a Zurich Instruments UHFLI 600 MHz lock-in amplifier for the probe tone. After frequency down-conversion of the microwave signal to the MHz range, the mechanical displacement excited by the probe tone is detected by the lock-in amplifier as a voltage signal. This signal is then fed back to the input of a AMD Spartan-7 FPGA (XC7S100) to create a time delay $\tau$. It is then added with the the masked data (mask $m(t)$ onto the input data $u(t)$) for the output of the FPGA. An AD835 analog multiplier is used to multiply the incident pump signal with the FPGA output, in order to modulate the pump force applied to the double-drum resonator.
Training: The displacement vector, created from the force amplitude modulation, will be linearly combined and give the final output $y(t)$ based on the formula:
@@
The $Z$ is the data displacement matrix consisting of $x(t)$ and $\lambda$ is the regularization parameter. When the training of the weight vector $W$ ends, we will test the performance of our system for different inputs [9, 11]. All computing processes are based on 2000 input bits. The first 100 inputs are discarded to eliminate offset effects from previous measurements. The subsequent 1500 inputs are used for training, and the remaining 400 inputs are used for prediction.
@@
X.Z. conceived the experiments. L.F. and X.Z. fabricated the samples in the clean room. P.L. and X.Z. developed the experimental setup. T.F., L.F., and X.Z. conducted the experiments and analyzed the results. T.F. and X.Z. wrote the manuscript, L.F.and B.D-R reviewed the manuscript, G-E.C., J-Y.O., Y.T., Y.P. contributed to the discussions. X.Z. supervised the project and acquired the research funding.